OpenStudy (anonymous):

factoring trinomials 3p^2-2p-5

OpenStudy (anonymous):

It is (3p-5)(p+1)

OpenStudy (anonymous):

to check to see if it is right, just foil it back out 3p*p=3p^2 3p*1=3p -5*p=-5p -5*1=-5 so put it together to look like 3p^2-2p-5 :)